# Ultralight dark matter detection with trapped-ion interferometry

### v1

@@ Abstract
$\delta_{s}=\sqrt{\frac{2}{\omega\mu\sigma_{0}}}\ ,$ (S20)
@@
$\eta\equiv\frac{|\bm{B}_{\rm shield}|}{|\bm{B}_{\rm naive}|}\ ,$ (S21)
@@
$\eta^{-1}$ $=\Bigg{|}\frac{1}{3R^{3}\gamma^{3}\mu}\left(f_{s}(R,\Delta,\gamma,\mu)\sinh(\gamma\Delta)+f_{c}(R,\Delta,\gamma,\mu)\cosh(\gamma\Delta)\right)\Bigg{|}\ ,$ (S22)
$f_{s}(R,\Delta,\gamma,\mu)$ $=(R\gamma)^{4}-2R^{3}\Delta\gamma^{4}+(R\gamma)^{2}(1+2\mu^{2}+\gamma^{2}\Delta^{2})+R\Delta\gamma^{2}(\mu(3-2\mu)-1)-(\mu-1)(1+2\mu+\gamma^{2}\Delta^{2})\ ,$ (S23)
$f_{c}(R,\Delta,\gamma,\mu)$ $=3\mu(R\gamma)^{3}+R^{2}\Delta\gamma^{3}(1-4\mu)+R\Delta^{2}\gamma^{3}(\mu-1)+\gamma\Delta(2\mu^{2}-\mu-1)\ .$ (S24)
We use this unintuitive but fully general expression to produce the modulus of the shielding factors shown in Fig. S1, and to represent in Fig. 1 the effects on our expected sensitivity of two different shield choices.
@@
$\eta_{\gamma\Delta\to 0,\Delta\to 0}$ $\simeq 1-\frac{2\Delta}{3R}\frac{(\mu-1)^{2}}{\mu}\ ,$ (S25)
$\eta_{\gamma\Delta\to\infty,\Delta\to 0}$ $\simeq\Bigg{|}\left(\cosh(\gamma\Delta)+\gamma\Delta\left(\frac{R}{3\Delta\mu}-\frac{2}{3\mu}\right)\sinh(\gamma\Delta)\right)^{-1}\Bigg{|}\ .$ (S26)
The first limit is valid as long as $\mu\ll R/\Delta$. These expressions reproduce correctly the small- and large-mass shielding efficiencies shown in Fig. S1.
Figure S1: Shielding efficiency $\eta$ as defined in Section S.I.3 for various possible shields. We assume a spherical shield made of either copper ($\mu=1,~{}\sigma=2.55\times 10^{7}$ S/m) or stainless steel ($\mu=10^{3},~{}\sigma=2\times 10^{4}$ S/m), and take various configurations for the typical size $R$ and the thickness of the shield $\Delta$.
S.II Signal Analysis
In this section, we provide additional details regarding the analysis of time-dependent signals in trapped-ion interferometers. In particular, we show how to perform the analysis in frequency space, and how to take into account finite-sampling effects.
S.II.1 Conventions
@@
$\widetilde{f}(\omega)=\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}f(t)\,,\quad f(t)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}e^{-i\omega t}\widetilde{f}(\omega)\,.$ (S27)
@@
$P_{f}=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}\mathrm{d}t\,\langle|f(t)|^{2}\rangle=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}S_{f}(\omega)\,,$ (S28)
@@
$2\pi\delta(\omega-\omega^{\prime})S_{f}(\omega)=\langle\widetilde{f}(\omega)\widetilde{f}^{*}(\omega^{\prime})\rangle\,.$ (S29)
S.II.2 Phase shift signal
@@
$\Delta\Phi(t)$ $\simeq 2e\int_{t}^{t+\Delta t}\bm{B}\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t^{\prime}}\mathrm{d}t^{\prime}$ (S30)
@@
where the window function is defined in terms of Heaviside functions as $W(t^{\prime},t,\Delta t)=\Theta(t^{\prime}-t)\Theta(t+\Delta t-t^{\prime})$. The last line of Eq. (S30) is a convolution of the window function with the projection of the magnetic field component along $\hat{\bm{\Sigma}}$. Therefore, by the convolution theorem, the Fourier transform (FT) of Eq. (S30) is given by
$\widetilde{\Delta\Phi}(\omega)=2e\,\widetilde{W}(\omega,\Delta t)\widetilde{\bm{B}}(\omega)\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t}\,,$ (S31)
@@
$S_{\Delta\Phi}(\omega)$ $=8\pi e^{2}\,\left<\left|\widetilde{\bm{B}}(\omega)\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t}\right|^{2}\right>\left|\widetilde{W}(\omega,\Delta t)\right|^{2}\,$ (S32)
@@
S.II.3 Dark matter signal
@@
In light of these observations, we assume two different orientations for ALP and DP searches. In the former case, we assume a trap with $\hat{\bm{\Sigma}}\parallel\bm{B_{a}}$. The ALP-sourced magnetic field strength is maximal at approximately $(169^{\circ}\mathrm{W},18^{\circ}\mathrm{S})$, close to the Bikini Atoll in the Pacific Ocean, as shown in Fig. S2. At this location, $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0.22$ G. Close to the geomagnetic poles, $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0$ G. Note that this does not imply that the Earth’s magnetic field is zero at the geomagnetic poles, where instead it is maximal and $\gtrsim 0.6$ G. Simply, it states that the component of the magnetic field sourced by the ALP in the $\bm{\Phi}_{\ell m}$ direction is approximately zero. In our projection, we assume that the sensor is located in Los Angeles, where we anticipate $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0.11$ G. This geographical choice degrades the amplitude of the signal by a factor of two, a correction which, however, is still well within the error bars of our noise model. For DP searches, without loss of generality, we assume a sensor with $\hat{\bm{\Sigma}}\parallel\hat{\bm{\phi}}$.
Figure S2: Normalized magnitude of the magnetic field sourced by ALP dark matter, $\bm{B_{a}}$, in nT at the Earth’s surface. This is equivalent to taking $\mathbb{R}\left[-ig_{a\gamma\gamma}a\right](m_{a}R_{\oplus})=1$ and $r=R_{\oplus}$ in Eq. (S19).
For our choice of sensor orientation, one can compute the expected PSD of the DM signal for a given DM mass and coupling to photons. For the ALP case, it trivially follows that the DM-dependent part of the PSD in Eq. (S32) is given by
$\left|\widetilde{\bm{B}}_{a}(\omega)\cdot\hat{\bm{\Sigma}}\right|^{2}\right>=\left<\left|\widetilde{\bm{B}}_{a}(\omega)\right|^{2}\right>$ $\approx g_{a\gamma\gamma}^{2}\left\langle|a|^{2}\right\rangle(m_{a}R_{\oplus})^{2}(0.11~{}\text{G})^{2}$ (S33)
$=2g_{a\gamma\gamma}^{2}R_{\oplus}^{2}\rho_{{}_{\mathrm{DM}}}(0.11~{}\text{G})^{2}\,,$
where we used the definitions of $\bm{\Phi}_{\ell m}$ provided in Table S1, assumed a sensor located in Los Angeles and employed the statistical properties of the ALP field, namely $\langle|a|^{2}\rangle=2\rho_{{}_{\mathrm{DM}}}/m_{a}^{2}$. For the DP case, this same quantity is given by
$\left|\widetilde{\bm{B}}_{{}_{A^{\prime}}}(\omega)\cdot\hat{\bm{\Sigma}}\right|^{2}\right>$ $=\frac{1}{4}\epsilon^{2}m_{A^{\prime}}^{4}R_{\oplus}^{2}\left[\left\langle|A^{\prime}_{0}|^{2}\right\rangle\sin^{2}\theta+\frac{1}{2}\left\langle|A^{\prime}_{1}|^{2}\right\rangle\cos^{2}\theta+\frac{1}{2}\left\langle|A^{\prime}_{-1}|^{2}\right\rangle\cos^{2}\theta\right]$ (S34)
@@
S.II.4 Likelihood and test statistics
@@
$\mathcal{L}(d|\mathcal{M},\bm{\theta})=\prod_{k=1}^{N-1}\frac{\exp\left(-\frac{|\widetilde{d}^{(k)}|^{2}}{S_{s}^{(k)}+S_{n}^{(k)}}\right)}{\pi\left(S_{s}^{(k)}+S_{n}^{(k)}\right)}\,.$ (S35)
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})=-2\ln\left(\frac{\mathcal{L}(d|\mathcal{M},\{m_{{}_{\mathrm{DM}}},\kappa\},\hat{\hat{\bm{\theta}}}_{n})}{\mathcal{L}(d|\mathcal{M},\{m_{{}_{\mathrm{DM}}},\widehat{\kappa^{2}}\},\hat{\bm{\theta}}_{n})}\right)\Theta\left(\kappa^{2}-\widehat{\kappa^{2}}\right)\,,$ (S36)
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})$ $\approx\sum_{k}\left(\frac{S_{s}^{(k)}}{S_{n}^{(k)}}\right)^{2}$ (S37)
@@
$S_{s}(\omega)\simeq\frac{2\pi P_{{}_{\mathrm{DM}}}}{\Delta\omega_{{}_{\mathrm{DM}}}}\Theta\left(\omega-m_{{}_{\mathrm{DM}}}\right)\Theta\left(\Delta\omega_{{}_{\mathrm{DM}}}+m_{{}_{\mathrm{DM}}}-\omega\right)\,,$ (S38)
@@
$\mathrm{SNR}\simeq\frac{P_{{}_{\mathrm{DM}}}}{S_{n}(m_{{}_{\mathrm{DM}}})}\sqrt{T_{\mathrm{int}}\tau_{c}}\,.$ (S39)
Since $P_{{}_{\mathrm{DM}}}=\langle|\Delta\Phi_{{}_{\mathrm{DM}}}|^{2}\rangle\propto\kappa^{2}$, where $\langle|\Delta\Phi_{{}_{\mathrm{DM}}}|^{2}\rangle$ is the amplitude squared of the time-averaged DM-induced phase shift, the 95% upper limit on $\kappa$ scales with $(T_{\mathrm{int}}\tau_{c})^{-1/4}$ and is determined by solving Eq. (S39) for $\mathrm{SNR}\approx 1.64$.
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})\approx 2\ln\left(\frac{S_{s}^{(k)}}{S_{n}^{(k)}}\right)-2\,,$ (S40)
@@
$S_{s}^{(k)}=\frac{1}{\Delta\omega}\int_{\omega_{k}-\Delta\omega/2}^{\omega_{k}+\Delta\omega/2}\mathrm{d}\omega\,S_{s}(\omega)\simeq\frac{2\pi P_{{}_{\mathrm{DM}}}}{\Delta\omega}\,.$ (S41)
@@
$\mathrm{SNR}\simeq\frac{P_{{}_{\mathrm{DM}}}}{S_{n}(m_{{}_{\mathrm{DM}}})}T_{\mathrm{int}}\,,$ (S42)
the 95% upper limit on $\kappa$ scales with $(T_{\mathrm{int}})^{-1/2}$ and is determined by solving Eq. (S42) for $\mathrm{SNR}\approx 12.5$.
S.II.5 Finite sampling effects
@@
$\widetilde{\Delta\Phi}_{\mathrm{sampled}}(\omega)=\widetilde{\Delta\Phi}(\omega)+\sum_{k\neq 0}^{\infty}\widetilde{\Delta\Phi}(\omega+k\omega_{\star})\,,$ (S43)
@@
$S_{{\Delta\Phi},\mathrm{sampled}}(\omega)=S_{\Delta\Phi}(\omega)+\sum_{k\neq 0}^{\infty}S_{\Delta\Phi}(\omega+k\omega_{\star})\,,$ (S44)

### v2

@@ B.3 Shielding around the ion trap
$\delta_{s}=\sqrt{\frac{2}{\omega\mu\sigma_{0}}}\ ,$ (57)
@@
$\eta\equiv\frac{|\bm{B}_{\rm shield}|}{|\bm{B}_{\rm naive}|}\ ,$ (58)
@@
$\eta^{-1}$ $=\Bigg|\frac{1}{3R^{3}\gamma^{3}\mu}\left(f_{s}(R,\Delta,\gamma,\mu)\sinh(\gamma\Delta)+f_{c}(R,\Delta,\gamma,\mu)\cosh(\gamma\Delta)\right)\Bigg|\ ,$ (59)
with
$f_{s}(R,\Delta,\gamma,\mu)$ $=(R\gamma)^{4}-2R^{3}\Delta\gamma^{4}+(R\gamma)^{2}(1+2\mu^{2}+\gamma^{2}\Delta^{2})+R\Delta\gamma^{2}(\mu(3-2\mu)-1)$
$\qquad-(\mu-1)(1+2\mu+\gamma^{2}\Delta^{2})\ ,$ (60)
$f_{c}(R,\Delta,\gamma,\mu)$ $=3\mu(R\gamma)^{3}+R^{2}\Delta\gamma^{3}(1-4\mu)+R\Delta^{2}\gamma^{3}(\mu-1)+\gamma\Delta(2\mu^{2}-\mu-1)\ .$ (61)
We use this unintuitive but fully general expression to produce the modulus of the shielding factors shown in Fig. 3, and to represent in Fig. 1 the effects on our expected sensitivity of two different shield choices.
@@
$\eta_{\gamma\Delta\to 0,\Delta\to 0}$ $\simeq 1-\frac{2\Delta}{3R}\frac{(\mu-1)^{2}}{\mu}\ ,$ (62)
$\eta_{\gamma\Delta\to\infty,\Delta\to 0}$ $\simeq\Bigg|\Bigg[\cosh(\gamma\Delta)$
$+\gamma\Delta\left(\frac{R}{3\Delta\mu}-\frac{2}{3\mu}\right)\sinh(\gamma\Delta)\Bigg]^{-1}\Bigg|\ .$ (63)
The first limit is valid as long as $\mu\ll R/\Delta$. These expressions reproduce correctly the small- and large-mass shielding efficiencies shown in Fig. 3.
Figure 3: Shielding efficiency $\eta$ as defined in Appendix B.3 for various possible shields. We assume a spherical shield made of either copper ($\mu=1,~\sigma=2.55\times 10^{7}$ S/m) or stainless steel ($\mu=10^{3},~\sigma=2\times 10^{4}$ S/m), and take various configurations for the typical size $R$ and the thickness of the shield $\Delta$.
Appendix C Signal Analysis
In this appendix, we provide additional details regarding the analysis of time-dependent signals in trapped-ion interferometers. In particular, we show how to perform the analysis in frequency space, and how to take into account finite-sampling effects.
C.1 Conventions
@@
$\widetilde{f}(\omega)=\int_{-\infty}^{\infty}\mathrm{d}te^{i\omega t}f(t)\,,\quad f(t)=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}e^{-i\omega t}\widetilde{f}(\omega)\,.$ (64)
@@
$P_{f}=\lim_{T\rightarrow\infty}\frac{1}{2T}\int_{-T}^{T}\mathrm{d}t\,\langle|f(t)|^{2}\rangle=\int_{-\infty}^{\infty}\frac{\mathrm{d}\omega}{2\pi}S_{f}(\omega)\,$, (65)
@@
$2\pi\delta(\omega-\omega^{\prime})S_{f}(\omega)=\langle\widetilde{f}(\omega)\widetilde{f}^{*}(\omega^{\prime})\rangle\,.$ (66)
C.2 Phase shift signal
@@
$\Delta\Phi(t)$ $\simeq 2e\int_{t}^{t+\Delta t}\bm{B}\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t^{\prime}}\mathrm{d}t^{\prime}$ (67)
@@
where the window function is defined in terms of Heaviside functions as $W(t^{\prime},t,\Delta t)=\Theta(t^{\prime}-t)\Theta(t+\Delta t-t^{\prime})$. The last line of Eq. (67) is a convolution of the window function with the projection of the magnetic field component along $\hat{\bm{\Sigma}}$. Therefore, by the convolution theorem, the Fourier transform (FT) of Eq. (67) is given by
$\widetilde{\Delta\Phi}(\omega)=2e\,\widetilde{W}(\omega,\Delta t)\widetilde{\bm{B}}(\omega)\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t}\,$, (68)
@@
$S_{\Delta\Phi}(\omega)$ $=8\pi e^{2}\,\left<\left|\widetilde{\bm{B}}(\omega)\cdot\frac{\mathrm{d}\bm{\Sigma}}{\mathrm{d}t}\right|^{2}\right>\left|\widetilde{W}(\omega,\Delta t)\right|^{2}\,$ (69)
@@
C.3 Dark matter signal
@@
In light of these observations, we assume two different orientations for ALP and DP searches. In the former case, we assume a trap with $\hat{\bm{\Sigma}}\parallel\bm{B_{a}}$. The ALP-sourced magnetic field strength is maximal at approximately $(169^{\circ}\mathrm{W},18^{\circ}\mathrm{S})$, close to the Bikini Atoll in the Pacific Ocean, as shown in Fig. 4. At this location, $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0.22$ G. Close to the geomagnetic poles, $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0$ G. Note that this does not imply that the Earth’s magnetic field is zero at the geomagnetic poles, where instead it is maximal and $\gtrsim 0.6$ G. Simply, it states that the component of the magnetic field sourced by the ALP in the $\bm{\Phi}_{\ell m}$ direction is approximately zero. In our projection, we assume that the sensor is located in Los Angeles, where we anticipate $B_{a}/(g_{a\gamma\gamma}am_{a}R_{\oplus})\approx 0.11$ G. This geographical choice degrades the amplitude of the signal by a factor of two, a correction which, however, is still well within the error bars of our noise model. For DP searches, without loss of generality, we assume a sensor with $\hat{\bm{\Sigma}}\parallel\hat{\bm{\phi}}$.
Figure 4: Normalized magnitude of the magnetic field sourced by ALP dark matter, $\bm{B_{a}}$, in nT at the Earth’s surface. This is equivalent to taking $\mathbb{R}\left[-ig_{a\gamma\gamma}a\right](m_{a}R_{\oplus})=1$ and $r=R_{\oplus}$ in Eq. (56).
For our choice of sensor orientation, one can compute the expected PSD of the DM signal for a given DM mass and coupling to photons. For the ALP case, it trivially follows that the DM-dependent part of the PSD in Eq. (69) is given by
[…]
where we used the definitions of $\bm{\Phi}_{\ell m}$ provided in Table 1, assumed a sensor located in Los Angeles and employed the statistical properties of the ALP field, namely $\langle|a|^{2}\rangle=2\rho_{{}_{\mathrm{DM}}}/m_{a}^{2}$. For the DP case, this same quantity is given by
$\left<\left|\widetilde{\bm{B}}_{{}_{A^{\prime}}}(\omega)\cdot\hat{\bm{\Sigma}}\right|^{2}\right$> $=\frac{1}{4}\epsilon^{2}m_{A^{\prime}}^{4}R_{\oplus}^{2}\left[\left\langle|A^{\prime}_{0}|^{2}\right\rangle\sin^{2}\theta+\frac{1}{2}\left\langle|A^{\prime}_{1}|^{2}\right\rangle\cos^{2}\theta+\frac{1}{2}\left\langle|A^{\prime}_{-1}|^{2}\right\rangle\cos^{2}\theta\right]$ (71)
@@
C.4 Likelihood and test statistics
@@
$\mathcal{L}(d|\mathcal{M},\bm{\theta})=\prod_{k=1}^{N-1}\frac{\exp\left(-\frac{|\widetilde{d}^{(k)}|^{2}}{S_{s}^{(k)}+S_{n}^{(k)}}\right)}{\pi\left(S_{s}^{(k)}+S_{n}^{(k)}\right)}\,.$ (72)
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})$ $=-2\ln\left(\frac{\mathcal{L}(d|\mathcal{M},\{m_{{}_{\mathrm{DM}}},\kappa\},\hat{\hat{\bm{\theta}}}_{n})}{\mathcal{L}(d|\mathcal{M},\{m_{{}_{\mathrm{DM}}},\widehat{\kappa^{2}}\},\hat{\bm{\theta}}_{n})}\right)$ (73)
[…]
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})$ $\approx\sum_{k}\left(\frac{S_{s}^{(k)}}{S_{n}^{(k)}}\right)^{2}$ (74)
@@
$S_{s}(\omega)\simeq\frac{2\pi P_{{}_{\mathrm{DM}}}}{\Delta\omega_{{}_{\mathrm{DM}}}}\Theta\left(\omega-m_{{}_{\mathrm{DM}}}\right)\Theta\left(\Delta\omega_{{}_{\mathrm{DM}}}+m_{{}_{\mathrm{DM}}}-\omega\right)\,$, (75)
@@
$\mathrm{SNR}\simeq\frac{P_{{}_{\mathrm{DM}}}}{S_{n}(m_{{}_{\mathrm{DM}}})}\sqrt{T_{\mathrm{int}}\tau_{c}}\,.$ (76)
Since $P_{{}_{\mathrm{DM}}}=\langle|\Delta\Phi_{{}_{\mathrm{DM}}}|^{2}\rangle\propto\kappa^{2}$, where $\langle|\Delta\Phi_{{}_{\mathrm{DM}}}|^{2}\rangle$ is the amplitude squared of the time-averaged DM-induced phase shift, the 95% upper limit on $\kappa$ scales with $(T_{\mathrm{int}}\tau_{c})^{-1/4}$ and is determined by solving Eq. (76) for $\mathrm{SNR}\approx 1.64$.
@@
$q(m_{{}_{\mathrm{DM}}},\kappa^{2})\approx 2\ln\left(\frac{S_{s}^{(k)}}{S_{n}^{(k)}}\right)-2\,$, (77)
@@
$S_{s}^{(k)}=\frac{1}{\Delta\omega}\int_{\omega_{k}-\Delta\omega/2}^{\omega_{k}+\Delta\omega/2}\mathrm{d}\omega\,S_{s}(\omega)\simeq\frac{2\pi P_{{}_{\mathrm{DM}}}}{\Delta\omega}\,.$ (78)
@@
$\mathrm{SNR}\simeq\frac{P_{{}_{\mathrm{DM}}}}{S_{n}(m_{{}_{\mathrm{DM}}})}T_{\mathrm{int}}\,$, (79)
the 95% upper limit on $\kappa$ scales with $(T_{\mathrm{int}})^{-1/2}$ and is determined by solving Eq. (79) for $\mathrm{SNR}\approx 12.5$.
C.5 Finite sampling effects
@@
$\widetilde{\Delta\Phi}_{\mathrm{sampled}}(\omega)=\widetilde{\Delta\Phi}(\omega)+\sum_{k\neq 0}^{\infty}\widetilde{\Delta\Phi}(\omega+k\omega_{\star})\,$, (80)
@@
$S_{{\Delta\Phi},\mathrm{sampled}}(\omega)=S_{\Delta\Phi}(\omega)+\sum_{k\neq 0}^{\infty}S_{\Delta\Phi}(\omega+k\omega_{\star})\,$, (81)